# The PLATO field selection process

### v1

@@ 3.7 Synergy with TESS
Figure 12: Distribution of observed TESS targets in LOPS2, ortographic projection in Galactic coordinates. Left panel: all 37 9103791037\,91037 910 CTL targets within LOPS2 observed in short cadence by TESS from sector 1 to 82 included (black points), superimposed on the LOPS2 field. The yellow circle is the TESS southern CVZ (up to 35 sectors), the red circle is roughly the northern boundary of Camera 4, i.e., the sky region where at least 4 TESS sectors are available. Right panel: same, but the TESS FFI coverage in terms of TESS sectors is color-coded over an HEALPix level-6 grid.
@@
Among the 519 704519704519\,704519 704 unique TESS targets (Stassun et al., 2018, 2019) observed at the regular 2-min cadence up to Sector 82 included (i.e., August 2024, end of Cycle 6), 37 9103791037\,91037 910 lie within the approximate footprint of LOPS2 with an average areal density spanning the 5-20 stars per $\textrm{deg}^{2}$ range (Fig. 12, left panel; the two most evident over-densities are centered on the LMC and NGC 2516). Among these, 55% were observed for at least three TESS sectors (not always contiguous) and 15% for at least 13 sectors, or one year cumulated (mostly within the $\beta\lesssim-78^{\circ}$ cap). Slightly more than 40% of the available TESS short-cadence targets are not included in the current version of the PIC [^21] Being included in the PIC does not automatically imply being scheduled for observations or being included in the P1-P2-P4-P5 sample, even though the vast majority of them will do (Montalto et al., 2021)., mostly because their stellar parameters fall outside our $(T_{\mathrm{eff}},R_{\star})$ parameter space (M21), or, to a lesser extent, because they do not meet our magnitude requirements. All those remaining targets of scientific interest left out can be requested by the community through the PLATO GO program, if they are technically feasible.
@@ 3.8 Synergy with CHEOPS
Figure 13: Overlap between CHEOPS/Ariel and LOPS2 (orthographic projection in Galactic coordinates). Left panel: in magenta, sky area forbidden to CHEOPS due to the Sun Exclusion Angle currently set at $120^{\circ}$ (corresponding to $\beta<-60^{\circ}$ for the southern ecliptic cap). P1 targets are plotted as black dots. Right panel: Ariel targets from the current MRS. Confirmed targets belonging to the Ariel Tier 1/2/3 (see text for details) are plotted as yellow, orange and red dots, respectively, while the TESS candidate planets are plotted as gray crosses (see text for details). The yellow circle is the TESS CVZ.
@@ 3.9 Synergy with Ariel
Ariel is an ESA M-class mission currently in development, with a planned launch to L2 in 2029 (Tinetti et al., 2018, 2021), i. e., with a significant temporal overlap with the PLATO nominal mission. Based on a 1-m class telescope equipped with two near-to-mid infrared low-resolution spectrographs (covering the 1.1-7.8 $\mu$m range) and three photometric channels also employed as fine guidance system (FGS; 0.5-1.1 $\mu$m), Ariel will monitor transiting planets through emission spectroscopy, transmission spectroscopy and phase curves, with repeated observations to gather the required SNR. Ariel targets will be observed through a four-“Tier” approach: Tier 1 targets will undergo an extensive low-resolution reconnaissance survey, Tier 2 targets will be selected for a deeper survey at higher resolution, and Tier 3 is made of a relatively small number of benchmark planets to be explored for variability (Edwards & Tinetti, 2022). Tier 4 will be dedicated to phase curves and bespoke observations of targets of special interest, which are expected to be identified e.g. in Tier 1 or by space-based missions and ground-based surveys prior to Ariel.
@@
The current Ariel target list, called Mission Reference Sample (MRS) is a living catalog, started by Edwards & Tinetti (2022) and regularly updated [^22] https://github.com/arielmission-space/Mission_Candidate_Sample. The latest version (2024-07-09) contains 722 confirmed transiting planets, and 2025 candidates from the TESS TOI/ExoFOP database (Guerrero et al., 2021). The majority of MRS targets (58%) are hot Jupiters. Also, nearly 90% of the “confirmed planets” sub-sample have transits deeper than 1 mmag, i. e., are easily detectable by TESS at high SNR (or in most cases, even from ground-based facilities). A cross-match with the LOPS2 footprint reveals that PLATO will observe 60 confirmed planets from the MRS (22 in Tier-1, 29 in Tier-2, nine in Tier-3) belonging to 49 host stars (Fig. 13, right panel).
@@
Given the selection criteria of the MRS, PLATO will be a provider of new, additional targets for Ariel, mainly of a small (but scientifically very interesting) population of long-period transiting giants not (yet) discovered by TESS, or discovered as mono-transit events only (Magliano et al., 2024). Moreover, the PLATO vs. Ariel synergy will be strong, for the already known and future targets, on the extremely accurate stellar parameters delivered by PLATO through asteroseismological analysis (and in particular, ages; Goupil et al. 2024), the in-depth characterization of stellar activity and rotational periods (Breton et al., 2024), the refinement of the planetary ephemeris for targets difficult to be followed up from the ground, the extension of the temporal baseline for TTV studies (Borsato et al., 2022), and the accurate planetary masses obtained by the PLATO follow-up program (Rauer et al., 2024). On the hot Jupiters belonging to the brighter tail of the MRS, PLATO will also be able to detect the planetary albedo and/or heat redistribution behavior through phase curves (Shporer, 2017; Singh et al., 2019), in a few cases possibly with color information thanks to the FCAM coverage.
@@ 4 Conclusions
Meanwhile, effort will be spent on building the more effective target lists and defining criteria on how to allocate the PLATO telemetry resources (imagettes, centroids, light curves) among the selected targets. The selected field meets the mission scientific requirements on samples P1-P2-P4-P5 as defined by SciRD. In the near future, the PSWT will implement a prioritization metric able to define a subset of most valuable stars for the ground-based follow-up, called prime sample. For planets hosted by this sample, the PLATO mission will eventually provide planet candidate confirmation and high-resolution spectroscopy, with planetary mass measurements. The prioritization process will also allow prioritization of tPIC targets for imagettes acquisition, sampling timing, etc. This process, to be documented in a third paper of this series, does not separately consider the different samples. The prioritization metric is evaluated regardless of whether a star belongs to the P1, P2 or P5 sample. For instance, a K3 subgiant at $V\lesssim 11$ and $R_{\star}\simeq 3$ $R_{\odot}$ can be included in P1 with a similar $\textrm{NSR}\simeq 45$ ppm with respect to a K3V star at $V\gtrsim 11$ and $R_{\star}\simeq 0.8$ $R_{\odot}$ belonging to P5, due to the magnitude constraint. However, detecting a true Earth analog during LOPS would be impossible in the former case, while perfectly within the reach [^23] As a figure of reference, a photometric precision of 80 ppm in one hour enables the robust detection of a true Earth analog around a quiet Sun twin with three transits (Rauer et al., 2024). of PLATO in the latter, so that this observation will be given higher priority.
@@ Acknowledgements.
We thank the referee for his valuable comments and suggestion. We thank Hans Deeg for his useful comments.
@@
This research has made use of the SIMBAD database (operated at CDS, Strasbourg, France; Wenger et al. 2000), TOPCAT and STILTS (Taylor, 2005, 2006). This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP; DOI:10.26134/ExoFOP5) website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.
Appendix A Is my favorite target within the LOPS2?
@@
A more accurate representation of LOPS2, also including the NCAM sub-regions, can be built by approximating the field on a level-9 HEALPix grid, and then converting the footprints into IVOA [^24] International Virtual Observatory Alliance, https://ivoa.net/-compliant Multi-Order Coverage (MOC) regions (Fernique et al. 2014), through the mocpy code [^25] https://cds-astro.github.io/mocpy/. The MOC regions for 6, 12, 18, 24 NCAMs and the underlying HEALPix grid will be made available as online supplementary material. We emphasize that the actual sky coverage could be slightly different due to several factors (the in-orbit optical performances, the relative co-alignment of the cameras, the pointing accuracy, etc.) so all these regions should be used with some caution especially when a target of interest is close the the external boundaries or to the inner gaps.
@@ Appendix B Additional tables and figures
Figure 14: Optimization of the PLATO field rotation angle: number of P1 (left panel) and P5 (right panel) targets as a function of the rotation angle, for hypothetical PLATO fields centered at the LOPN1 (blue points and lines) or LOPS2 (green points and lines) coordinates. The continuous and dashed horizontal lines mark, for reference, the value of P1/P5 counts at $\varphi=0^{\circ}$ and a ±$\pm 1\%$ deviation from it, respectively. See Section 2.2 for more details.
Figure 15: Optimization of the PLATO field rotation angle: PLATO test fields centered at the LOPS2 coordinates (Table 2) but rotated by $0^{\circ}$, $10^{\circ}$, $20^{\circ}$, $30^{\circ}$, $40^{\circ}$, $50^{\circ}$, $60^{\circ}$, $70^{\circ}$, $80^{\circ}$ (in reading order); projection is Galactic. The $\varphi=0^{\circ}$ rotation angle corresponds to the nominal LOPS2 geometry. As in Fig. 6, the brightest stars from the Yale Bright Star Catalog are plotted as circles of different colors: magenta ($V<1$), red ($21<V<$ 2), yellow ($32<V<$ 3). The TESS southern CVZ is marked with a yellow circle. See Section 2.2 for more details.
@@
Figure 17: All-sky Aitoff projection in Equatorial coordinates of LOPS2 and LOPN1, showing the formal constraints for the selection of the PLATO LOP fields and the synergies with other missions. The two pink circles represent the $|\beta|>63^{\circ}$ technical requirement for the center of the LOP fields (“allowed region”), implying that the overall envelopes of every allowed field choice are two ecliptic caps at $|\beta|\gtrsim 38^{\circ}$ (green circles). LOPS2 (lower left) and LOPN1 (upper right) are plotted with blue shades according to the number of co-pointing cameras, as in Fig. 1. The footprints of CoRoT (red), Kepler (magenta), and K2 (green) are over-plotted together with the TESS continuous viewing zone at $|\beta|\gtrsim 78^{\circ}$ (yellow circle). The background gray layer is color-coded according to the areal density of $G<13.5$ stars from Gaia DR3. The celestial equator and poles are marked with a red line and crosses, respectively.
Figure 18: All-sky Aitoff projection in Ecliptic coordinates of LOPS2 and LOPN1, showing the formal constraints for the selection of the PLATO LOP fields and the synergies with other missions. The two pink circles represent the $|\beta|>63^{\circ}$ technical requirement for the center of the LOP fields (“allowed region”), implying that the overall envelopes of every allowed field choice are two ecliptic caps at $|\beta|\gtrsim 38^{\circ}$ (green circles). LOPS2 (lower left) and LOPN1 (upper right) are plotted with blue shades according to the number of co-pointing cameras, as in Fig. 1. The footprints of CoRoT (red), Kepler (magenta), and K2 (green) are over-plotted together with the TESS continuous viewing zone at $|\beta|\gtrsim 78^{\circ}$ (yellow circle). The background gray layer is color-coded according to the areal density of $G<13.5$ stars from Gaia DR3. The celestial equator and poles are marked with a red line and crosses, respectively.
@@
Figure 22: Upper panels: histograms of the expected orbital period of a planet at the center of the habitable zone according to the the Kasting et al. (1993) definition ($a_{\mathrm{hz}}\textrm{ [au]}=\sqrt{L_{\star}/L_{\odot}}$) and Kepler’s laws, for the P1+P2 (left side) and P5 (right side) samples within the LOPS2 field (see Section 3.1 for details). The Sun/Earth value (1 yr) is marked with dashed vertical line as reference. Both the distribution of the whole sample (in gray) and for a few subsamples of interest is shown: main-sequence F stars (blue), main-sequence G and K stars (green) and FGK subgiants (orange). The definition of these sub-samples in terms of $T_{\mathrm{eff}}$ and $R_{\star}$ can be found in N22, Section 6.3. Lower panels: same, but for a planet located at the inner edge of the habitable zone according to the the Kasting et al. (1993) definition ($a_{\mathrm{hz}}\textrm{ [au]}=0.77\times\sqrt{L_{\star}/L_{\odot}}$).
Figure 23: Spatial distribution of the P1-P2-P4-P5 targets (in reading order) over the LOPS2 field. P1, P4 and P5 targets are represented as color-coded density maps based on an HEALPix level-5, while the sparse and almost perfectly isotropic distribution of the P2 sample is plotted with individual red points. See also Section 3.1 for a discussion of the sample properties.

### v3

@@ 3.7 Synergy with TESS
Figure 12: Distribution of observed TESS targets in LOPS2, orthographic projection in galactic coordinates. Left panel: All 37 9103791037\,91037 910 CTL targets within LOPS2 observed in short cadence by TESS from sector 1 to 82 included (black points), superimposed on the LOPS2 field. The yellow circle is the TESS southern CVZ (up to 35 sectors), the red circle is roughly the northern boundary of Camera 4, that is, the sky region where at least 4 TESS sectors are available. Right panel: Same but the TESS FFI coverage in terms of TESS sectors is color coded over an HEALPix level-6 grid.
@@
Among the 519 704519704519\,704519 704 unique TESS targets (Stassun et al., 2018, 2019) observed at the regular 2-min cadence up to Sector 82 included (i.e., August 2024, end of Cycle 6), 37 9103791037\,91037 910 lie within the approximate footprint of LOPS2 with an average areal density spanning the 5-20 stars per $\textrm{deg}^{2}$ range (Fig. 12, left panel; the two most evident over-densities are centered on the LMC and NGC 2516). Among these, 55% were observed for at least three TESS sectors (not always contiguous) and 15% for at least 13 sectors, or one year cumulated (mostly within the $\beta\lesssim-78^{\circ}$ cap). Slightly more than 40% of the available TESS short-cadence targets are not included in the current version of the PIC, [^21] Being included in the PIC does not automatically imply being scheduled for observations or being included in the P1-P2-P4-P5 sample, even though the vast majority of them will do (Montalto et al., 2021). mostly because their stellar parameters fall outside our $(T_{\mathrm{eff}},R_{\star})$ parameter space (M21), or, to a lesser extent, because they do not meet our magnitude requirements. All those remaining targets of scientific interest left out can be requested by the community through the PLATO GO program, if they are technically feasible.
@@ 3.8 Synergy with CHEOPS
Figure 13: Overlap between CHEOPS/Ariel and LOPS2 (orthographic projection in galactic coordinates). Left panel: Sky area forbidden to CHEOPS (in magenta) due to the Sun Exclusion Angle currently set at $120^{\circ}$ (corresponding to $\beta<-60^{\circ}$ for the southern ecliptic cap). P1 targets are plotted as black dots. Right panel: Ariel targets from the current MRS. Confirmed targets belonging to the Ariel Tier 1/2/3 (see text for details) are plotted as yellow, orange and red dots, respectively, while the TESS candidate planets are plotted as gray crosses (see text for details). The yellow circle is the TESS CVZ.
@@ 3.9 Synergy with Ariel
Ariel is an ESA M-class mission currently in development, with a planned launch to L2 in 2029 (Tinetti et al., 2018, 2021), i. e., with a significant temporal overlap with the PLATO nominal mission. Based on a 1-m class telescope equipped with two near-to-mid infrared low-resolution spectrographs (covering the 1.1-7.8 $\mu$m range) and three photometric channels also employed as fine guidance system (FGS; 0.5-1.1 $\mu$m), Ariel will monitor transiting planets through emission spectroscopy, transmission spectroscopy and phase curves, with repeated observations to gather the required SNR. Ariel targets will be observed through a four-“Tier” approach: Tier 1 targets will undergo an extensive low-resolution reconnaissance survey, Tier 2 targets will be selected for a deeper survey at higher resolution, and Tier 3 is made of a relatively small number of benchmark planets to be explored for variability (Edwards & Tinetti, 2022). Tier 4 will be dedicated to phase curves and bespoke observations of targets of special interest, which are expected to be identified, for example, in Tier 1 or by space-based missions and ground-based surveys prior to Ariel.
@@
The current Ariel target list, called Mission Reference Sample (MRS) is a living catalog, started by Edwards & Tinetti (2022) and regularly updated. [^22] https://github.com/arielmission-space/Mission_Candidate_Sample The latest version (2024-07-09) contains 722 confirmed transiting planets, and 2025 candidates from the TESS TOI/ExoFOP database (Guerrero et al., 2021). The majority of MRS targets (58%) are hot Jupiters. Also, nearly 90% of the “confirmed planets” sub-sample have transits deeper than 1 mmag, i. e., are easily detectable by TESS at high SNR (or in most cases, even from ground-based facilities). A cross-match with the LOPS2 footprint reveals that PLATO will observe 60 confirmed planets from the MRS (22 in Tier-1, 29 in Tier-2, nine in Tier-3) belonging to 49 host stars (Fig. 13, right panel).
@@
Given the selection criteria of the MRS, PLATO will be a provider of new, additional targets for Ariel, mainly of a small (but scientifically very interesting) population of long-period transiting giants not (yet) discovered by TESS, or discovered as mono-transit events only (Magliano et al., 2024). Moreover, the PLATO versus Ariel synergy will be strong, for the already known and future targets, on the extremely accurate stellar parameters delivered by PLATO through asteroseismological analysis (and in particular, ages; Goupil et al. 2024), the in-depth characterization of stellar activity and rotational periods (Breton et al., 2024), the refinement of the planetary ephemeris for targets difficult to be followed up from the ground, the extension of the temporal baseline for TTV studies (Borsato et al., 2022), and the accurate planetary masses obtained by the PLATO follow-up program (Rauer et al., 2024). On the hot Jupiters belonging to the brighter tail of the MRS, PLATO will also be able to detect the planetary albedo and/or heat redistribution behavior through phase curves (Shporer, 2017; Singh et al., 2019), in a few cases possibly with color information thanks to the FCAM coverage.
@@ 4 Conclusions
Meanwhile, effort will be spent on building the more effective target lists and defining criteria on how to allocate the PLATO telemetry resources (imagettes, centroids, light curves) among the selected targets. The selected field meets the mission scientific requirements on samples P1-P2-P4-P5 as defined by SciRD. In the near future, the PSWT will implement avprioritization metric able to define a subset of most valuable stars for the ground-based follow-up, called prime sample. For planets hosted by this sample, the PLATO mission will eventually provide planet candidate confirmation and high-resolution spectroscopy, with planetary mass measurements. The prioritization process will also allow prioritization of tPIC targets for imagettes acquisition, sampling timing, etc. This process, to be documented in a third paper of this series, does not separately consider the different samples. The prioritization metric is evaluated regardless of whether a star belongs to the P1, P2 or P5 sample. For instance, a K3 subgiant at $V\lesssim 11$ and $R_{\star}\simeq 3$ $R_{\odot}$ can be included in P1 with a similar $\textrm{NSR}\simeq 45$ ppm with respect to a K3V star at $V\gtrsim 11$ and $R_{\star}\simeq 0.8$ $R_{\odot}$ belonging to P5, due to the magnitude constraint. However, detecting a true Earth analog during LOPS would be impossible in the former case, while perfectly within the reach [^23] As a figure of reference, a photometric precision of 80 ppm in one hour enables the robust detection of a true Earth analog around a quiet Sun twin with three transits (Rauer et al., 2024). of PLATO in the latter, so that this observation will be given higher priority.
@@
Data availability
The table with the confirmed transiting planets in LOPS2, the table with the TESS candidates in LOPS2 and the MOC regions of LOPS2 are publicly available on Zenodo: zenodo14720127.
@@ Acknowledgements.
We thank the referee for his valuable comments and suggestions. We thank Hans Deeg for his useful comments.
@@
This research has made use of the SIMBAD database (operated at CDS, Strasbourg, France; Wenger et al. 2000), TOPCAT and STILTS (Taylor, 2005, 2006), UCAC4-RPM (Nascimbeni et al., 2016), TEPCAT (Southworth, 2011). This research has made use of the Exoplanet Follow-up Observation Program (ExoFOP; DOI:10.26134/ExoFOP5) website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.
Appendix A The LOPS2 footprint
@@
A more accurate representation of LOPS2, also including the NCAM sub-regions, can be built by approximating the field on a level-9 HEALPix grid, and then converting the footprints into IVOA [^24] International Virtual Observatory Alliance, https://ivoa.net/-compliant Multi-Order Coverage (MOC) regions (Fernique et al. 2014), through the mocpy code [^25] https://cds-astro.github.io/mocpy/. The MOC regions for 6, 12, 18, 24 NCAMs and the underlying HEALPix grid are available online262626zenodo14720127. We emphasize that the actual sky coverage could be slightly different due to several factors (the in-orbit optical performances, the relative co-alignment of the cameras, the pointing accuracy, etc.) so all these regions should be used with some caution especially when a target of interest is close the the external boundaries or to the inner gaps.
@@ Appendix B Additional tables and figures
Figure 14: Optimization of the PLATO field rotation angle. The number of P1 (left panel) and P5 (right panel) targets are shown as a function of the rotation angle, for hypothetical PLATO fields centered at the LOPN1 (blue points and lines) or LOPS2 (green points and lines) coordinates. The continuous and dashed horizontal lines mark, for reference, the value of P1/P5 counts at $\varphi=0^{\circ}$ and a ±$\pm 1\%$ deviation from it, respectively (see Section 2.2 for more details).
Figure 15: Optimization of the PLATO field rotation angle. The PLATO test fields are centered at the LOPS2 coordinates (Table 2) but rotated by $0^{\circ}$, $10^{\circ}$, $20^{\circ}$, $30^{\circ}$, $40^{\circ}$, $50^{\circ}$, $60^{\circ}$, $70^{\circ}$, $80^{\circ}$ (in reading order); projection is galactic. The $\varphi=0^{\circ}$ rotation angle corresponds to the nominal LOPS2 geometry. As in Fig. 6, the brightest stars from the Yale Bright Star Catalog are plotted as circles of different colors: magenta ($V<1$), red ($21<V<$ 2), yellow ($32<V<$ 3). The TESS southern CVZ is marked with a yellow circle (see Section 2.2 for more details).
@@
Figure 17: All-sky Aitoff projection in Equatorial coordinates of LOPS2 and LOPN1, showing the formal constraints for the selection of the PLATO LOP fields and the synergies with other missions. The two pink circles represent the $|\beta|>63^{\circ}$ technical requirement for the center of the LOP fields (“allowed region”), implying that the overall envelopes of every allowed field choice are two ecliptic caps at $|\beta|\gtrsim 38^{\circ}$ (green circles). LOPS2 (lower left) and LOPN1 (upper right) are plotted with blue shades according to the number of co-pointing cameras, as in Fig. 1. The footprints of CoRoT (red), Kepler (magenta), and K2 (green) are over-plotted together with the TESS continuous viewing zone at $|\beta|\gtrsim 78^{\circ}$ (yellow circle). The background gray layer is color coded according to the areal density of $G<13.5$ stars from Gaia DR3. The celestial equator and poles are marked with a red line and crosses, respectively.
Figure 18: All-sky Aitoff projection in Ecliptic coordinates of LOPS2 and LOPN1, showing the formal constraints for the selection of the PLATO LOP fields and the synergies with other missions. The two pink circles represent the $|\beta|>63^{\circ}$ technical requirement for the center of the LOP fields (“allowed region”), implying that the overall envelopes of every allowed field choice are two ecliptic caps at $|\beta|\gtrsim 38^{\circ}$ (green circles). LOPS2 (lower left) and LOPN1 (upper right) are plotted with blue shades according to the number of co-pointing cameras, as in Fig. 1. The footprints of CoRoT (red), Kepler (magenta), and K2 (green) are over-plotted together with the TESS continuous viewing zone at $|\beta|\gtrsim 78^{\circ}$ (yellow circle). The background gray layer is color coded according to the areal density of $G<13.5$ stars from Gaia DR3. The celestial equator and poles are marked with a red line and crosses, respectively.
@@
Figure 22: Orbital periods corresponding to the habitable zone around PLATO targets. Upper panels: Histograms of the expected orbital period of a planet at the center of the habitable zone according to the the Kasting et al. (1993) definition ($a_{\mathrm{hz}}\textrm{ [au]}=\sqrt{L_{\star}/L_{\odot}}$) and Kepler’s laws, for the P1+P2 (left side) and P5 (right side) samples within the LOPS2 field (see Section 3.1 for details). The Sun/Earth value (1 yr) is marked with dashed vertical line as reference. Both the distribution of the whole sample (in gray) and for a few subsamples of interest is shown: main-sequence F stars (blue), main-sequence G and K stars (green) and FGK subgiants (orange). The definition of these sub-samples in terms of $T_{\mathrm{eff}}$ and $R_{\star}$ can be found in N22, Section 6.3. Lower panels: Same but for a planet located at the inner edge of the habitable zone according to the the Kasting et al. (1993) definition ($a_{\mathrm{hz}}\textrm{ [au]}=0.77\times\sqrt{L_{\star}/L_{\odot}}$).
Figure 23: Spatial distribution of the P1-P2-P4-P5 targets (in reading order) over the LOPS2 field. P1, P4 and P5 targets are represented as color coded density maps based on an HEALPix level-5, while the sparse and almost perfectly isotropic distribution of the P2 sample is plotted with individual red points. See also Section 3.1 for a discussion of the sample properties.